Provide complete solution
Let {€1, єг, єз, є4} be a basis for a vector space V of dinnension 4. Define a linear transformation on V such that 4) 2.
Answer #1


Solution: Given linear transformation A: V V satisfies where {єд4-1 is a basis for V. Let x є V and x-(x1+x2, x3, x4)- Then 4-1 xeï Thus x = (z1,22, x2, x3) E ker(A) → x = (-T2-x3, 12, 23, 0)=T2(-1, 1, 0, 0)+13(-1,0,1,0) € span(m= (-1, 1,0, 0),O2= (-1,0,1,0)) Inn(A) = {A(x) : x equation 3) 2. E V} = span(q,g) and dim(Inn(A)) = 2, (by 3. Note that α1 =-€1 + €2 (X1,X2, X3, X4) € ker(A)nInn(A) =÷ x-(x1,22 , 0, 0) where x1+T2=0 x = (-T2,T2.0, 0) = x2(-1,1,0,0) = x201 → ker(A)n 1m(A) = span fa1)